# Lithium Niobate Photonic Topological Insulator-based Multi-Wavelength Optical Demultiplexer with Piezoelectric Switch-Off

### v1

@@ 1 Introduction
Lithium niobate (LN) is a popular synthetic crystal for photonic and telecommunication applications. It has a broad optical transparency region spanning from the visible to the mid-infrared region, high refractive index [27], and favorable electro-optic properties [28, 29, 30]. These properties make LN an excellent candidate for integrated photonics [31, 32, 33]. The advancements in manufacturing wafer-scale, high-quality thin films of lithium niobate-on-insulator and breakthroughs in nanofabrication techniques have made LN nanophotonic components possible and capable of outperforming traditional photonic components[31]. In this paper, a lithium niobate photonic crystal has been used to design the valley PTI. While numerous works on silicon-based PTI have been done, lithium niobate has not yet been extensively demonstrated in this field, particularly in the telecommunication region. There has been an investigation on the rotated Weyl physics on nonlinear Lithium Niobate-on-Insulator Chips [34] and a report of lithium niobate valley photonic crystal waveguide [35]. However, in previous literature, there has been no demonstration of a valley photonic topological insulator on a lithium niobate platform that operates in the telecommunication region, which can also be tuned to a wide range of operating wavelengths with equal bandwidth in each case. Furthermore, previous works do not include demultiplexing operations and piezoelectric switching capabilities. The use of lithium niobate instead of the traditional silicon is motivated by two factors. The first factor is the wide transparency region of the LN material. The primary regions of operation shown in this work are the telecommunication regions of 1310 nm and 1550 nm. Using silicon for the PTI design with other operating wavelengths is difficult due to the high attenuation. However, the wide transparency region of lithium niobate allows the operating wavelength to be adjusted over a much more comprehensive range without significant attenuation. The second reason for using LN in our work is its prominent inverse piezoelectric effect [28, 36, 37, 38, 39]. The operation of valley PTIs strongly depends on the spatial arrangement and the lattice dimensions. An external voltage induces stress in the lattice structures due to the inverse piezoelectric effect of LN. We show that if the magnitude of the external voltage is above a certain threshold, the stress-induced distortions in the lattice can remove topological protection at the operating wavelength. Based on this principle, we propose a piezoelectric switch-off mechanism for the optical channel.
@@ 2.1 Simulation setup
All optical simulations were done by solving Maxwell’s equations using the finite-difference time-domain (FDTD) method [46]. The FDTD simulations were performed using the commercial software Lumerical FDTD: 3D Electromagnetic Simulator [47]. The first step was generating the structure in Fig. 1(a). We wrote our custom code using the Lumerical scripting language to develop it. Each unit of the lattice is shown in the inset of Fig. 1(d). The dimensions of the structure depend on the lattice constant, . The length of the equilateral triangles is for the photonic crystal lattice with a symmetry, and and for the photonic crystal lattice with a symmetry, the spacing between the triangles is 0.4, and the width of the channel is also 0.4. The structure consists of 82 units in the x direction and 40 units in the y direction, 20 units for each of the upper and lower photonic crystals. The dispersive refractive indices of the LN slab are shown in Fig. 1(b) [27]. A TE mode source was used as input. Simulations were done with a mesh override region surrounding each triangle for accurate convergence. This is necessary for high resolution around the sharp corners and edges of the triangles. The maximum mesh step for the mesh override regions was chosen based on the smallest wavelength and the smallest dimension of the simulation. A maximum mesh step of 24 nm in the x, y, and z directions was taken for structures with 850 nm, while a smaller mesh step of 6 nm was taken for structures with = 850 nm and smaller because of the smaller triangles. A separate scripting code generated the structure with the sharp turns in Fig. 3(c). The defects in Fig. 3(e) were introduced by manually removing air holes near the channel. All post-processing was done in MATLAB.
@@ 2.3 Transmittance spectrum, spatial transmittance profile and normalized spatial power distribution
The source was placed at the start of the channel. A 2-dimensional (2D) frequency domain power monitor along the YZ plane was used to calculate the spatial transmittance profile in Fig. 1(c). A 1-dimensional (1D) frequency domain power monitor placed at the end of the channel with vertical dimensions equal to the channel width was used to calculate the transmittance spectrum and the spatial transmittance profile of the channel shown in Fig. 2, 3, 4, 5 and 6. A similar monitor placed at the start of the channel was used to calculate the backscattering from the channel shown in Fig. 3. A 2-dimensional (2D) frequency domain power monitor was placed in the XY plane across the structure to calculate the normalized spatial power distributions shown in Fig. 2, 3, 4, 5 and 6. For the results in Fig. 4(a), we generated structures by sweeping the lattice constant from 600 nm to 1000 nm and analyzed the peaks from the transmittance spectrum in each case. Both 2D and 3D FDTD simulations were required.
@@ 2.4 Piezoelectric simulations
The simulation of the inverse piezoelectric effect of LN was done using the finite element method (FEM) [49]. At first, the entire structure was generated in the simulation environment. Ground was set to the lower boundary, and the input potential was set to the upper boundary to generate the potential distribution shown in Fig. 6(a). The step potential in Fig. 6(b) was used as input. The strain-charge relation of the piezoelectric material, along with transient simulations, was done to generate the spatial stress profile in Fig. 6(c) and 6(h) and the magnitude of stress as a function of time in Fig. 6(d). The necessary equations are shown in Section 3.3. The transmittance spectrum in Fig. 6(e) and the normalized spatial power distribution in Fig. 6(f) and 6(g) of the stress-induced structure were then calculated by FDTD simulations.
@@ 3.1 Multi-wavelength photonic topological insulator
Our main design is shown in Fig. 1(a) with the predicted pathway of light. It consists of a lithium niobate slab with a honeycomb lattice. Each unit cell has a length equal to the lattice constant, , and contains two inverted equilateral triangles, with lengths and , as shown in the inset of Fig. 1(d). The periodic arrangement of these lattices in the slab forms two different photonic crystal structures with opposite spatial orientations. The positions of the large and small triangles are interchanged in the upper and lower photonic crystals. This forms a valley PTI with the interface between the two structures acting as the unidirectional pathway. The width of the channel is taken as 0.4 to allow proper light propagation. The dispersive refractive index of the LN slab is shown in Fig. 1(b) [27]. Light propagates only in the XY plane. It is confined in the z direction by total internal reflection, as shown by the spatial transmittance profile in Fig. 1(c).
Fig. 1(d) shows the photonic band structure of our design. The inset of Fig. 1(d) shows the Brillouin Zone path, the unit cell of the photonic crystal slab, and the dimensions of the unit cell. We found a Dirac cone at the high symmetry point in the momentum space for the photonic crystal lattice with a symmetry, having . Introducing asymmetry by setting and breaks the inversion symmetry of the lattice and reduces the lattice symmetry to . This lifts the degeneracy at the point and opens up a band gap, as seen by the red plot in Fig. 1(d). Similar behaviors in the band structure were observed in previous works on Si PTI [50, 51]. The parameters and were chosen because they open a well-defined band gap in the band structure at the desired location. The thickness of the slab, , will depend on the availability of wafers. It is better to keep small to suppress higher-order slab modes in the out-of-plane direction. In Fig. 1(d), the frequency in the y-axis has been normalized. The location of the band gap and the Dirac cone are given in this normalized unit. The value of is adjusted based on the desired operating wavelength. The anisotropic nature of the LN crystal resulted in a directional band gap. This will cause imperfect insulation of the bulk. However, we present spatial transmittance profiles and the transmittance through the bulk that verify that the transmission is confined to the channel and the scattering into the bulk is negligible.
@@
Through Fig. 2, we study the existence of the topologically protected edge state, wavelength selectivity, channel confinement, and bulk insulation of our design. Fig. 2(a) is the projected band structure, calculated with the periodic section of the main structure in the inset. This section contains the interface in the middle, surrounded by 20 unit cells along each side of the y-axis. The unit cells above and below have opposite spatial orientations. The edge state (the blue curve) crosses the band gap (shaded green region) between the bulk bands (shaded grey regions). This edge state is approximately linear in the band gap, indicating a constant group velocity. The dashed black line represents the operating wavelength in the center of the band gap. Our initial target wavelengths are the telecommunication regions of 1310 nm and 1550 nm. The structures for = 766 nm and = 904 nm were considered for achieving these operating wavelengths. The values of the lattice constant, , are chosen so that the ratio / is at the center of the band gap. For these two particular cases, / = 766 nm/1310 nm = 904 nm/1550 nm 0.583. Fig. 2(b) is the transmittance spectrum of the channel, calculated using the FDTD method. The vertical dimension of the monitor was equal to the width of the channel, so the spectrum shows only the components that propagate to the end of the channel. As we can see, high transmittance is observed around our desired wavelength while other components are suppressed. Similar shapes can be seen in the spectrum of previous PTI designs [50, 52]. The wavelength range of the topological edge states is localized around the operating wavelength. Spectral components outside the range of the edge mode cannot transmit through the channel. That is why we observe a bell-shaped transmittance with the spectrum diminishing as it moves further from the operating wavelength. Fig. 2(c) and 2(d) show the normalized spatial power distribution across the XY plane of the entire structure at = 1310 nm and = 1550 nm for = 766 nm and = 904 nm, respectively. In both cases, the optical signal is well confined within the channel and insulated from the bulk. No bulk scattering can be observed at the wavelengths of interest. To further verify that the operating wavelengths do not scatter into the bulk, we present the spatial transmittance profiles in Fig. 2(e) and 2(f). Each has been calculated with monitors at the end of the channel and plotted against the y-direction coordinates at operating wavelengths corresponding to their respective lattice constant. The operating wavelength remains confined to the channel. In Fig. 2(g) and 2(h), we compare the spectral components transmitted through the channel and the spectral components scattered into the bulk. Fig. 2(g) shows that the components around = 1310 nm do not scatter into the bulk for the structure with = 766 nm, and Fig. 2(h) shows that the components around = 1550 nm do not scatter into the bulk for the structure with = 904 nm. The absence of bulk scattering verifies the presence of the band gap at the operating wavelength, and the edge transmission of the operating wavelength quantitatively verifies the existence of an edge state.
@@
Equation 2 gives the value of the lattice constant, , for the operating wavelength. The other parameters depend on , so they will change accordingly. To further substantiate this claim, we consider two structures for = 850 nm and = 950 nm and compare the transmittance with the previous results. Fig. 4(b) shows that the transmittance spectrum taken from the end of the channel is similar for the four cases, but the peaks have shifted. Fig. 4(c) and 4(d) are the normalized spatial power distribution across the structures for = 850 nm and = 950 nm, respectively, taken at their peak wavelength. They show that the optical power is confined to the channel and insulated at the bulk despite the change in the lattice dimensions. So, our design can operate as a photonic topological insulator for a wide range of wavelengths. The large transparency window, high refractive index, and low material loss of LN permit this multi-wavelength feature and open up the possibility of using our design at numerous operating wavelengths.
@@ 3.2 Demonstration as an optical demultiplexer
An important application of the wavelength-selective nature of the transmittance spectrum of our PTI design is to isolate wavelength components, while the bulk insulation is useful for channel isolation and, hence, low crosstalk. Fig. 5 presents the performance of our design as a 2-channel optical demultiplexer. We demonstrate operation at 1310 nm and 1550 nm. Fig. 5(a) shows the demultiplexer structure. Two structures with = 766 nm and = 904 nm are generated and stacked side by side. The resulting structure has two channels or interfaces with two input and two output ports. A y-splitter waveguide made of LN is used to couple the optical source into the input ports of the channel. LN is chosen as the waveguide material to maximize mode matching and minimize coupling loss at the interface of the waveguide and the slab for higher coupling efficiency. Fig. 5(b) and 5(c) demonstrate the normalized spatial power distribution across the structure at = 1310 nm and = 1550 nm, respectively. The y-splitter transmits both components. But the lower channel supports only the = 1310 nm component, and the upper channel supports only the = 1550 nm component. The topological protection forces the optical power to be insulated at the bulk and, hence, the two channels to be isolated. Fig. 5(d) and 5(e) are the plots of the transmittance spectra in dB of both channels taken within selective regions. The transmittance spectra show that each channel has high transmittance around its operating wavelength with good bandwidth while suppressing unwanted components. The crosstalk in the lower channel goes as low as 54 dB, and in the upper channel, it goes as low as 27 dB. The wavelengths of 1310 nm and 1550 nm have been used as the carrier wavelengths. However, using lithium niobate for the PTI allows for a wider selection of PTI operating wavelengths, hence, the demultiplexer carrier wavelengths. A demultiplexer structure for any two of the operating wavelengths mapped in Fig. 4(a) can be formed using the same design. All we require is a change in the lattice constant.
@@ 3.3 Piezoelectric switch-off
where , , and are tensors that represent the induced strain, applied stress field, and applied electric field, respectively. is the elastic compliance tensor at a constant electric field, and is the transpose of the piezoelectric strain coefficient tensor. As shown in the inset of Fig. 6(b), the electric field is along axis 3, and no external stress field has been applied. Taking these factors into consideration and utilizing the crystal symmetry of LN, the strain components of interest on the plane of the photonic crystal structure can be expressed by
@@
These equations model the inverse piezoelectric effect on the photonic crystal due to the applied voltage. The electric field due to the 2 V external voltage induces constant stress in the 1-3 plane of the structure, as shown in Fig. 6(c). Fig. 6(d) shows the magnitude of the stress as a function of time. The magnitude settles to 0.3 MPa. The inverse piezoelectric strain acts along the 1 and 3 axes, as given by the components in Eq. (4). The boundary transmission of this strained structure is studied using FDTD simulations. The results are presented in Fig. 6(e), 6(f), and 6(g). The transmittance spectrum of the channel with the external voltage turned on and off is compared in Fig. 6(e). The wavelength range of = 1530 nm to = 1565 nm has been used considering the operating wavelength of the = 904 nm structure. When the applied voltage is turned off, the unstrained structure shows a transmittance peak at the operating wavelength. After a 2 V external voltage is applied, the transmittance curve of the strained structure shows that the operating wavelength decreases to less than 5 of its original value. A minimum of 2 V potential is required for the transmittance to be considered off. A higher voltage will further decrease the transmittance. However, a higher voltage will also induce a higher stress. A high stress across the structure might cause permanent defects in the photonic crystal. Considering these factors, the magnitude of the applied voltage and the generated stress are kept within typical reported values [36, 38, 39]. Applying the voltage parallel to the plane of the photonic crystal structure instead of along the height of the slab allows strain components and , parallel to the photonic crystal plane, to have higher values at acceptable voltages. Fig. 6(f) and 6(g) show the normalized spatial power distribution across the XY plane of the structure at the operating wavelength. The results are for when the applied potential is respectively 0 V and 2 V. The high transmission through the channel and topological protection observed in the unstrained structure, as shown in Fig. 6(f), decreases significantly in the strained structure after the potential is applied, as shown in Fig. 6(g). Fig. 6(h) shows the effect of fabrication defects on the stress distribution. The defects are modeled as missing air gaps near the channel, indicated by the dotted red circles in the figure. According to Eq. (3), (4), and (5), the inverse piezoelectric response depends on the voltage distribution and the electric field due to the distribution. If the number of defects does not significantly alter the voltage distribution, the stress distribution will remain the same. Fig. 6(g) demonstrates that introducing defects does not change the stress distribution, and the settled value remains 0.3 MPa. Thus, a minimum of 2 V applied potential removes topological protection and turns off transmission through the channel. Removing the applied voltage removes the induced strain and restores transmission. These results introduce a piezoelectric method to switch off the topological properties and transmission through the channel. Although the results for the structure with = 904 nm were presented, the mechanism also applies to structures with other lattice constants. Each channel of the demultiplexer can be similarly switched by this piezoelectric method.

### v3

@@ 1 Introduction
Lithium niobate (LN) is a popular synthetic crystal for photonic and telecommunication applications. It has a broad optical transparency region spanning from the visible to the mid-infrared region, high refractive index [27], and favorable electro-optic properties [28, 29, 30]. These properties make LN an excellent candidate for integrated photonics [31, 32, 33]. The advancements in manufacturing wafer-scale, high-quality thin films of lithium niobate-on-insulator and breakthroughs in nanofabrication techniques have made LN nanophotonic components possible and capable of outperforming traditional photonic components[31]. In this paper, a lithium niobate photonic crystal has been used to design the valley PTI. While numerous works on silicon-based PTI have been done, lithium niobate has not yet been extensively demonstrated in this field, particularly in the telecommunication region. There has been an investigation on the rotated Weyl physics on nonlinear Lithium Niobate-on-Insulator Chips [34] and a report of lithium niobate valley photonic crystal waveguide [35]. However, in previous literature, there has been no demonstration of a valley photonic topological insulator on a lithium niobate platform that operates in the telecommunication region, which can also be tuned to a wide range of operating wavelengths with equal bandwidth in each case. Furthermore, previous works do not include demultiplexing operations and piezoelectric switching capabilities. The use of lithium niobate instead of the traditional silicon is motivated by two factors. The first factor is the wide transparency region of the LN material. The primary regions of operation shown in this work are the telecommunication regions of 1310 nm and 1550 nm. Using silicon for the PTI design with other operating wavelengths is difficult due to the high attenuation. However, the wide transparency region of lithium niobate allows the operating wavelength to be adjusted over a much more comprehensive range without significant attenuation. The second reason for using LN in our work is its prominent inverse piezoelectric effect [36, 37, 38, 39]. The operation of valley PTIs strongly depends on the spatial arrangement and the lattice dimensions. An external voltage induces stress in the lattice structures due to the inverse piezoelectric effect of LN. We show that if the magnitude of the external voltage is above a certain threshold, the stress-induced distortions in the lattice can remove topological protection at the operating wavelength. Based on this principle, we propose a piezoelectric switch-off mechanism for the optical channel.
@@ 2.1 Simulation setup
All optical simulations were done by solving Maxwell’s equations using the finite-difference time-domain (FDTD) method [46]. The FDTD simulations were performed using the commercial software Lumerical FDTD: 3D Electromagnetic Simulator [47]. The first step was generating the structure in Fig. 1(a). We wrote our custom code using the Lumerical scripting language to develop it. Each unit of the lattice is shown in Fig. 1(a). The dimensions of the structure depend on the lattice constant, . The length of the equilateral triangles is for the photonic crystal lattice with a symmetry, and and for the photonic crystal lattice with a symmetry, the spacing between the triangles is 0.4, and the width of the channel is also 0.4. The structure consists of 82 units in the x direction and 40 units in the y direction, 20 units for each of the upper and lower photonic crystals. The dispersive refractive indices of the LN slab are shown in Fig. 1(b) [27]. A TE mode source was used as input. Simulations were done with a mesh override region surrounding each triangle for accurate convergence. This is necessary for high resolution around the sharp corners and edges of the triangles. The maximum mesh step for the mesh override regions was chosen based on the smallest wavelength and the smallest dimension of the simulation. A maximum mesh step of 24 nm in the x, y, and z directions was taken for structures with 850 nm, while a smaller mesh step of 6 nm was taken for structures with 850 nm because of the smaller triangles. A separate scripting code generated the structure with the sharp turns. Defects were introduced by manually removing air holes near the channel. All post-processing was done in MATLAB.
@@ 2.3 Transmittance spectrum, spatial transmittance profile and normalized spatial power distribution
The source was placed at the start of the channel. A 2-dimensional (2D) frequency domain power monitor along the YZ plane was used to calculate the spatial transmittance profile in Fig. 1(c). A 1-dimensional (1D) frequency domain power monitor placed at the end of the channel with vertical dimensions equal to the channel width was used to calculate the transmittance spectrum and the spatial transmittance profile of the channel. A similar monitor placed at the start of the channel was used to calculate the backscattering from the channel. A 2-dimensional (2D) frequency domain power monitor was placed across the XY plane across the structure to calculate the normalized spatial power distributions. To find the relation between peak wavelength and lattice constant, we generated structures by sweeping the lattice constant from 600 nm to 1000 nm and analyzed the peaks from the transmittance spectrum in each case. Both 2D and 3D FDTD simulations were required.
@@ 2.4 Piezoelectric simulations
The simulation of the inverse piezoelectric effect of LN was done using the finite element method (FEM) [49]. At first, the entire structure was generated in the simulation environment. Ground was set to the lower boundary, and the input potential was set to the upper boundary to generate the potential distribution. A step potential was used as input. The strain-charge relation of the piezoelectric material, along with transient simulations, was done to generate the spatial stress profile and the magnitude of stress as a function of time. The necessary equations are shown in Section 3.3. The transmittance spectrum and the normalized spatial power distributions of the stress-induced structure were then calculated using FDTD simulations.
@@ 3.1 Multi-wavelength photonic topological insulator
Our main design is shown in Fig. 1(a) with the predicted pathway of light. It consists of a lithium niobate slab with a honeycomb lattice. Each unit cell has a length equal to the lattice constant, , and contains two inverted equilateral triangles, with lengths and , as shown in Fig. 1(a). The periodic arrangement of these lattices in the slab forms two different photonic crystal structures with opposite spatial orientations. The positions of the large and small triangles are interchanged in the upper and lower photonic crystals. This forms a valley PTI with the interface between the two structures acting as the unidirectional pathway. The width of the channel is taken as 0.4 to allow proper light propagation. The dispersive refractive index of the LN slab is shown in Fig. 1(b) [27]. Light propagates only in the XY plane. It is confined in the z direction by total internal reflection, as shown by the spatial transmittance profile in Fig. 1(c).
Fig. 1(d) shows the photonic band structure of our design. The inset of Fig. 1(d) shows the Brillouin Zone path. We found a Dirac cone at the high symmetry point in the momentum space for the photonic crystal lattice with a symmetry, having . Introducing asymmetry by setting and breaks the inversion symmetry of the lattice and reduces the lattice symmetry to . This lifts the degeneracy at the point and opens up a band gap, as seen by the red plot in Fig. 1(d). Similar behaviors in the band structure were observed in previous works on Si PTI [50, 51]. The parameters and were chosen because they open a well-defined band gap in the band structure at the desired location. The thickness of the slab, , will depend on the availability of wafers. It is better to keep small to suppress higher-order slab modes in the out-of-plane direction. In Fig. 1(d), the frequency in the y-axis has been normalized. The location of the band gap and the Dirac cone are given in this normalized unit. The value of is adjusted based on the desired operating wavelength. The anisotropic nature of the LN crystal resulted in a directional band gap. This will cause imperfect insulation of the bulk. However, we present spatial transmittance profiles and the transmittance through the bulk that verify that the transmission is confined to the channel and the scattering into the bulk is negligible.
@@
Through Fig. 2, we study the existence of the topologically protected edge state, wavelength selectivity, channel confinement, and bulk insulation of our design. Fig. 2(a) is the projected band structure showing the edge state (red line). The edge state is approximately linear in the band gap, indicating a constant group velocity. Our initial target wavelengths are the telecommunication regions of 1310 nm and 1550 nm. The structures for = 766 nm and = 904 nm were considered for achieving these operating wavelengths. The values of the lattice constant, , are chosen so that the ratio / is at the center of the band gap. For these two particular cases, / = 766 nm/1310 nm = 904 nm/1550 nm 0.583. Fig. 2(b) is the transmittance spectrum of the channel, calculated using the FDTD method. The vertical dimension of the monitor was equal to the width of the channel, so the spectrum shows only the components that propagate to the end of the channel. As we can see, high transmittance is observed around our desired wavelength while other components are suppressed. Similar shapes can be seen in the spectrum of previous PTI designs [50, 52]. The wavelength range of the topological edge states is localized around the operating wavelength. Spectral components outside the range of the edge mode cannot transmit through the channel. That is why we observe a bell-shaped transmittance with the spectrum diminishing as it moves further from the operating wavelength. Fig. 2(c) and 2(d) show the normalized spatial power distribution across the XY plane of the entire structure at = 1310 nm and = 1550 nm for = 766 nm and = 904 nm, respectively. In both cases, the optical signal is well confined within the channel and insulated from the bulk. No bulk scattering can be observed at the wavelengths of interest. To further verify that the operating wavelengths do not scatter into the bulk, we present the spatial transmittance profiles in Fig. 2(e) and 2(f). Each has been calculated with monitors at the end of the channel and plotted against the y-direction coordinates at operating wavelengths corresponding to their respective lattice constant. The operating wavelength remains confined to the channel. In Fig. 2(g) and 2(h), we compare the spectral components transmitted through the channel and the spectral components scattered into the bulk. Fig. 2(g) shows that the components around = 1310 nm do not scatter into the bulk for the structure with = 766 nm, and Fig. 2(h) shows that the components around = 1550 nm do not scatter into the bulk for the structure with = 904 nm. The absence of bulk scattering verifies the presence of the band gap at the operating wavelength, and the edge transmission of the operating wavelength quantitatively verifies the existence of an edge state.
@@
Eq. 2 gives the value of the lattice constant, , for the operating wavelength. The other parameters depend on , so they will change accordingly. To further substantiate this claim, we consider two structures for = 850 nm and = 950 nm and compare the transmittance with the previous results. Fig. 4(b) shows that the transmittance spectrum taken from the end of the channel is similar for the four cases, but the peaks have shifted. Fig. 4(c) and 4(d) are the normalized spatial power distribution across the structures for = 850 nm and = 950 nm, respectively, taken at their peak wavelength. They show that the optical power is confined to the channel and insulated at the bulk despite the change in the lattice dimensions. So, our design can operate as a photonic topological insulator for a wide range of wavelengths. The large transparency window, high refractive index, and low material loss of LN permit this multi-wavelength feature and open up the possibility of using our design at numerous operating wavelengths.
@@ 3.2 Demonstration as an optical demultiplexer
An important application of the wavelength-selective nature of the transmittance spectrum of our PTI design is to isolate wavelength components, while the bulk insulation is useful for channel isolation and, hence, low crosstalk. Fig. 5 presents the performance of our design as a 2-channel optical demultiplexer. We demonstrate operation at 1310 nm and 1550 nm. Fig. 5(a) shows the demultiplexer structure. Two structures with = 766 nm and = 904 nm are generated and stacked side by side. The resulting structure has two channels or interfaces with two input and two output ports. A y-splitter waveguide made of LN is used to couple the optical source into the channel input ports. LN is chosen as the waveguide material to maximize mode matching and minimize coupling loss at the interface of the waveguide and the slab for higher coupling efficiency. Fig. 5(b) and 5(c) demonstrate the normalized spatial power distribution across the structure at = 1310 nm and = 1550 nm, respectively. The y-splitter transmits both components. But the lower channel supports only the = 1310 nm component, and the upper channel supports only the = 1550 nm component. The topological protection forces the optical power to be insulated at the bulk and, hence, the two channels to be isolated. Fig. 5(d) and 5(e) are the plots of the transmittance spectra in dB of both channels taken within selective regions. The transmittance spectra show that each channel has high transmittance around its operating wavelength with good bandwidth while suppressing unwanted components. The crosstalk in the lower channel goes as low as 54 dB, and in the upper channel, it goes as low as 27 dB. The wavelengths of 1310 nm and 1550 nm have been used as the carrier wavelengths. However, using lithium niobate for the PTI allows for a wider selection of PTI operating wavelengths, hence, the demultiplexer carrier wavelengths. A demultiplexer structure for any two of the operating wavelengths mapped in Fig. 4(a) can be formed using the same design. All we require is a change in the lattice constant.
@@ 3.3 Piezoelectric switch-off
where , , and are tensors that represent the induced strain, applied stress field, and applied electric field, respectively. is the elastic compliance tensor at a constant electric field, and is the transpose of the piezoelectric strain coefficient tensor. As shown in the inset of Fig. 6(b), the electric field is along axis "3", and no external stress field has been applied. Taking these factors into consideration and utilizing the crystal symmetry of LN, the strain components of interest on the plane of the photonic crystal structure can be expressed by
@@
These equations model the inverse piezoelectric effect on the photonic crystal due to the applied voltage. The electric field due to the 2 V external voltage induces constant stress in the 1-3 plane of the structure, as shown in Fig. 6(c). Fig. 6(d) shows the magnitude of the stress as a function of time. The magnitude settles to 0.3 MPa. The inverse piezoelectric strain acts along the "1" and "3" axes, as given by the components in Eq. (4). The boundary transmission of this strained structure is studied using FDTD simulations. The results are presented in Fig. 6(e), 6(f), and 6(g). The transmittance spectrum of the channel with the external voltage turned on and off is compared in Fig. 6(e). The wavelength range of = 1530 nm to = 1565 nm has been used considering the operating wavelength of the = 904 nm structure. When the applied voltage is turned off, the unstrained structure shows a transmittance peak at the operating wavelength. After a 2 V external voltage is applied, the transmittance curve of the strained structure shows that the operating wavelength decreases to less than 5 of its original value. A minimum of 2 V potential is required for the transmittance to be considered off. A higher voltage will further decrease the transmittance. However, a higher voltage will also induce a higher stress. A high stress across the structure might cause permanent defects in the photonic crystal. Considering these factors, the magnitude of the applied voltage and the generated stress are kept within typical reported values [37, 38, 39]. Applying the voltage parallel to the plane of the photonic crystal structure instead of along the height of the slab allows strain components and , parallel to the photonic crystal plane, to have higher values at acceptable voltages. Fig. 6(f) and 6(g) show the normalized spatial power distribution across the XY plane of the structure at the operating wavelength. The results are for when the applied potential is respectively 0 V and 2 V. The high transmission through the channel and topological protection observed in the unstrained structure, as shown in Fig. 6(f), decreases significantly in the strained structure after the potential is applied, as shown in Fig. 6(g). Fig. 6(h) shows the effect of fabrication defects on the stress distribution. The defects are modeled as missing air gaps near the channel, indicated by the dotted red circles in the figure. According to Eq. (4) and (5), the inverse piezoelectric response depends on the voltage distribution and the electric field due to the distribution. If the number of defects does not significantly alter the voltage distribution, the stress distribution will remain the same. Fig. 6(g) demonstrates that introducing defects does not change the stress distribution, and the settled value remains 0.3 MPa. Thus, a minimum of 2 V applied potential removes topological protection and turns off transmission through the channel. Removing the applied voltage removes the induced strain and restores transmission. These results introduce a piezoelectric method to switch off the topological properties and transmission through the channel. Although the results for the structure with = 904 nm were presented, the mechanism also applies to structures with other lattice constants. Each channel of the demultiplexer can be similarly switched by this piezoelectric method.